# Controllable Generation with Text-to-Image Diffusion Models: A Survey

### v1

@@ 5.3 Weight Fusion
[…]
where $\{W_{n,l}^{k},W_{n,l}^{v}\}_{n=1}^{N}$ represent the corresponding updated key and value matrices for added $N$ concepts and $C_{\text{reg}}$ is a randomly sampled text features for regularization. The objective of Equation.13 is intuitively designed to ensure that the words in the target captions are consistently aligned with the values derived from the concept matrices that have undergone fine-tuning. Similarly, Mix-of-Show [181] introduces the gradient fusion, updating weight $W$ by $W=\arg\min_{W}\sum_{i=1}^{n}\left\|(W_{0}+\Delta W_{i})X_{i}-WX_{i}\right\|_{F}^{2}$ where $X_{i}$ represents the input activation of the $i$-th concept, and $|\cdot|_{F}$ denotes the Frobenius norm. To integrate subject-centric and style-centric conditions, ZipLoRA [182] merges LoRA-style weights by minimizing the difference between subject/style images generated by the mixed and original LoRA models and the cosine similarity between the columns of content and style LoRAs. Po et al. [183] present orthogonal adaption to replace LoRA in fine-tuning, encouraging the customized models to have orthogonal residual weights for efficient fusion.
@@ 5.4 Attention-based Integration
Attention-based integration methods modulate attention maps to strategically position subjects within the synthesized image, allowing for precise control over where and how each condition is represented in the final composition.
@@ 5.5 Guidance Composition
Guidance composition is an integration mechanism for synthesizing images under multiple conditions, integrating the independent denoising results of each condition. This process is mathematically represented as:
@@
To integrate multiple concepts, Decompose and Realign [185] obtains the corresponding $\mathcal{M}_{i}$ by their cross-attention map. Similarly, Face-diffuser [89] presents a saliency-adaptive noise fusion method to combine results from a text-driven diffusion model and a proposed subject-augmented diffusion model. Besides, Cao et al. [119] proposes the generalized classifier-free guidance (GCFG) for concept-centric personalization and integrates concept guidance and control guidance by manually setting intensities $w_{i}$.
@@ 6.1 Universal Conditional Score Prediction Framework
DiffBlender [186] is proposed to incorporate conditions from diverse types of modalities. It categorizes conditions into multiple types to employ different techniques for guiding generation. First, image-form conditions, which contain spatially rich information, are injected in ResNet Blocks [227]. Then, spatial conditions, including grounding box and keypoints, are passed through a local self-attention module to accurately locate the desired positions of synthesized results. Moreover, non-spatial conditions like color palette and style are concatenated with textual tokens through a global self-attention module and then fed into cross-attention layers. Additionally, Emu2 [187] leverages a large generative multimodal model with 37 billion parameters for task-agnostic in-context learning to construct a universal controllable T2I generation framework. After trained on a mix of high-quality datasets, it is capable of accepting a mixture of conditions like text, locations, and image as input, and generating images in context.
@@ 7 Applications
In this section, we focus on innovative methods that utilize novel conditions in the generation process to address specific tasks. By emphasizing these pioneering approaches, we aim to highlight how conditional generation is not only reshaping the landscape of content creation but also broadening the horizons of creativity and functionality in various fields. The subsequent discussions will provide insights into the transformative impact of these models and their potential in diverse applications.
@@ 7.1 Image Manipulation
Advancements in the control of pre-trained text-to-image diffusion models have allowed for more versatile image editing techniques. For instance, inspired by DreamBooth[42], SINE[234] constructs the text prompt for fine-tuning the pretrained text-to-image model by the source image as ”a photo/painting of a [***] [class]” and edits the image by a novel model-based classifier-free guidance. Moreover, the versatility of control conditions further enhances the editing process by integrating conditions beyond mere text. For example, Choi et al. [235] customize the diffusion model to employ specific elements from the reference image as editing criteria, such as substituting the cat in the source image with the cat appearance in the reference image. Additionally, several approaches [236] utilize spatial manipulation by spatial control, employing sketches or layouts to intuitively adjust the arrangement of elements within an image.
@@ 7.4 Text/Image-to-3D Generation
Text/image-to-3D task aims to reconstruct 3D representations from text descriptions or images (pairs). In the early days, the task developed slowly due to their reliance on expensive 3D annotations. The strong open-world knowledge of large-scale text-to-image diffusion models brings effective solutions without heavy 3D annotation requirements. For instance, by personalizing latent diffusion model[23], Zero-1-to-3 [243] builds a viewpoint-conditioned image translation diffusion model that generates multiple views of the input object image. Then the paired images are fed into a NeRF[244] model to do reconstruction.
Recent advancements in text/image-to-3D generation represent a significant milestone with the development of Score Distillation Sampling (SDS) loss. This innovative approach, introduced by DreamFusion[245], marks a successful adaptation of large-scale 2D diffusion models for 3D generation. Through SDS, the control method of the text-to-image model can be transferred to text-to-3D generation. Typically, DreamBooth3D[246] combines DreamBooth[42] and DreamFusion[245] that personalizes text-to-3d generative models from a few captured images of a subject. Similarly, some approaches[247, 248, 249, 248] adapt ControlNet[190] to the SDS process, enabling the control of 3D generation through spatial signals (e.g., depth map, sketch).

### v2

@@ 4.3 Weight Fusion
$\text{s.t. }WC^{T}$ $=V,\text{ where }C=[c_{1}\ldots c_{N}]^{T}$
$\text{ and }V$ $=[W_{1}c_{1}^{T}\ldots W_{N}c_{N}^{T}]^{T}$
where $\{W_{n,l}^{k},W_{n,l}^{v}\}_{n=1}^{N}$ represent the corresponding updated key and value matrices for added $N$ concepts and $C_{\text{reg}}$ is a randomly sampled text features for regularization. The objective of Eq. 18 is intuitively designed to ensure that the words in the target captions are consistently aligned with the values derived from the concept matrices that have undergone fine-tuning. Similarly, Mix-of-Show [75] introduces the gradient fusion, updating weight $W$ by $W=\arg\min_{W}\sum_{i=1}^{n}\left\|(W_{0}+\Delta W_{i})X_{i}-WX_{i}\right\|_{F}^{2}$ where $X_{i}$ represents the input activation of the $i$-th concept, and $|\cdot|_{F}$ denotes the Frobenius norm. To integrate subject-centric and style-centric conditions, ZipLoRA [231] merges LoRA-style weights by minimizing the difference between subject/style images generated by the mixed and original LoRA models and the cosine similarity between the columns of content and style LoRAs. Po et al. [198] present orthogonal adaption to replace LoRA in fine-tuning, encouraging the customized models to have orthogonal residual weights for efficient fusion.
@@ 4.4 Attention-based Integration
Attention-based integration methods modulate attention maps to strategically position subjects within the synthesized image, allowing for precise control over where and how each condition is represented in the final composition [163, 75, 131].
@@ 4.5 Guidance Composition
Guidance composition is an integration mechanism for synthesizing images under multiple conditions, integrating the independent denoising results of each condition [106, 274, 22, 269, 24, 270]. This process is mathematically represented as:
@@
To integrate multiple concepts, Decompose and Realign [269] obtains the corresponding $\mathcal{M}_{i}$ by their cross-attention map. Similarly, Face-diffuser [274] presents a saliency-adaptive noise fusion method to combine results from a text-driven diffusion model and a proposed subject-augmented diffusion model.
Besides, to realize controllable generation in user-specific domain, Cao et al. [22] train a null-text UNet to provide domain guidance and utilize the original diffusion prior to provide control guidance.
@@ 5.1 Universal Conditional Score Prediction Framework
DiffBlender [123] is proposed to incorporate conditions from diverse types of modalities. It categorizes conditions into multiple types to employ different techniques for guiding generation. First, image-form conditions, which contain spatially rich information, are injected in ResNet Blocks [86]. Then, spatial conditions, including grounding box and keypoints, are passed through a local self-attention module to accurately locate the desired positions of synthesized results. Moreover, non-spatial conditions like color palette and style are concatenated with textual tokens through a global self-attention module and then fed into cross-attention layers. Additionally, Emu2 [247] leverages a large generative multimodal model with 37 billion parameters for task-agnostic in-context learning to construct a universal controllable T2I generation framework.
@@ 5.2 Universal Condition-Guided Score Estimation
Figure 8: Illustration of the application of controllable text-to-image generation. The condition is marked in blue background. Examples are sourced from [331, 244, 296, 252, 278, 157, 85].
@@ 6 Applications
In this section, we focus on innovative methods that utilize novel conditions in the generation process to address specific tasks. By emphasizing these pioneering approaches, we aim to highlight how conditional generation is not only reshaping the landscape of content creation but also broadening the horizons of creativity and functionality in various fields. The subsequent discussions will provide insights into the transformative impact of these models and their potential in diverse applications. We illustrate the example of the applications in Fig. 8.
@@ 6.1 Image Manipulation
Advancements in the control of pre-trained text-to-image diffusion models have allowed for more versatile image editing techniques. For instance, inspired by DreamBooth[221], SINE[330] constructs the text prompt for fine-tuning the pretrained text-to-image model by the source image as ”a photo/painting of a [*] [class]” and edits the image by a novel adapter-based classifier-free guidance. Moreover, the versatility of control conditions further enhances the editing process by integrating conditions beyond mere text. For example, Choi et al. [48] customize the diffusion model to employ specific elements from the reference image as editing criteria, such as substituting the cat in the source image with the cat’s appearance in the reference image. Recently, Zhou et al. [345] modify the score estimation in multi-turn editing, introducing a dual-objective Linear Quadratic Regulators (LQR)to effectively mitigate error accumulation.
@@ 6.4 Text/Image-to-3D Generation
Text/image-to-3D task aims to reconstruct 3D representations from text descriptions or images (pairs). Recent advancements in text/image-to-3D generation represent a significant milestone with the development of Score Distillation Sampling (SDS) loss. This innovative approach, introduced by DreamFusion[200], marks a successful adaptation of large-scale 2D diffusion models for 3D generation. Through SDS, the control method of the text-to-image model can be transferred to text-to-3D generation. Typically, DreamBooth3D[210] combines DreamBooth[221] and DreamFusion[200] that personalizes text-to-3d generative models from a few captured images of a subject. Similarly, some approaches[40, 105, 309, 105] adapt ControlNet[323] to the SDS process, enabling the control of 3D generation through spatial signals (e.g., depth map, sketch).
6.5 World Model
The condition-injection mechanisms provide effective support for the development of video-generation-based world models. A representative example is camera-controlled video generation [280, 85, 9, 311, 8, 218, 10], which focuses on aligning sequences of camera parameters with the diffusion-based video generation process. For instance, ReCamMaster [10] incorporates camera parameters into the original DiT blocks via frame-dimension conditioning. Similarly, AC3D [8] introduces camera information through lightweight 128-dimensional DiT-SX blocks. In addition, several approaches tackle this problem in a training-free manner [95, 151, 238]. Typically, WorldForge [238], which leverages 3D/4D foundation models [267] to project video frames into a static point cloud. The point cloud is then adjusted according to camera trajectories and subsequently used as guidance for video generation.
@@ 7 Conclusion
Although controllable generation with text-to-image diffusion models has achieved remarkable progress, several promising directions remain open for future exploration:
(1) Towards a universal and cross-modal control paradigm. Existing controllable diffusion models are often tailored to specific conditions or tasks. Future research could focus on developing unified and generalizable control frameworks capable of flexibly accommodating diverse forms of conditions, including spatial, semantic, and multimodal inputs, within a single generative system. Extending controllability beyond text-to-image synthesis to other modalities such as audio, video, and 3D generation will further enhance the model’s adaptability and cross-modal reasoning ability, paving the way toward general multimodal intelligence.
(2) Building world models with controlling mechanisms. The condition-injection and controllability principles of diffusion models provide a strong foundation for constructing world models based on video generation. Future studies could explore how diffusion-based systems simulate dynamic, camera-controllable environments while maintaining temporal–spatial consistency. Such world models will play a crucial role in connecting generative modeling, embodied AI, and interactive virtual environments.